# ViViDex: Learning Vision-based Dexterous Manipulation from Human Videos

### v1

@@ III-C The Vision-based Policy
Though our state-based policy can work well for different manipulation tasks, it requires robot proprioceptive states and object states as inputs. However, reliably estimating object states is often not trivial in practice. To alleviate this issue, we propose a vision-based policy that only takes robot states and 3D scene point clouds as inputs. To generate training data for the visual policy, we rollout the optimized state-based policy and generate diverse physically plausible trajectories. During the rollout process, we compute the 3D point clouds ${{\rm PC_{w}}\in\mathbb{R}^{N\times 3}}$ from three depth cameras in the simulator, where $\rm PC_{w}$ is represented under the world coordinate system (i.e., the center of the table) and $N$ is the number of points.
Figure 4: Motion retargeting results for selected DexYCB [43] videos. The three rows show different initial object poses for each object category.
As shown in Figure 3, inspired by [5], we propose to transform $\rm PC_{w}$ into the desired target coordinate system $\rm PC_{t}$, which make the model more aware of the target position for control predictions. To capture dense interaction features between the robot and the object, we additionally transform $\rm PC_{w}$ into six different robot joint coordinate systems (i.e., one wrist and five fingers). Finally, we combine point clouds representations under different coordinate systems ${\rm PC}\in\mathbb{R}^{N\times 24}$ and feed them into PointNet [67]. Our visual policy predicts control commands based on extracted visual features and robot states. We train our visual policy using behavior cloning from generated physically plausible trajectories.
@@ IV Experiments
We conduct extensive experiments to evaluate our state-based and visual policies on three manipulation tasks, namely relocate, pour and place inside tasks.
@@ IV-A Experimental Setting
Video dataset. The DexYCB [43] dataset contains human demonstration videos of hand-object interactions for 20 YCB objects with 20 videos per object. Following [23], we focus on five objects including mustard bottle, tomato soup can, sugar box, large clamp and mug. Due to the unavailability of DexMV videos [23], we select videos from the DexYCB dataset that shares high visual similarities with DexMV videos. For each object, we choose one video which is most similar to those used in DexMV and use this video for most of our experiments; the top row in Figure 4 presents motion retargeting results for these videos. For the final experiment, we add additional initial configurations of the objects to the test set (different from [23]) and use two additional videos per object with different initial object positions for training, as illustrated in the last two rows of Figure 4. We use hand and object poses provided by the dataset to perform motion retargeting and generate reference trajectories.
Simulation benchmark. We follow previous work [23, 27] and use the MuJoCo physics simulator [68] together with the Adroit robot hand [19]. The robot hand consists of 30 degrees of freedom (DoF). We set the control frequency and simulation frequency as 40HZ and 400HZ, respectively. As discussed above, we use the same testing configuration as in [23] except for the final experiment, where we use a more difficult setting, and evaluate the performance of our method on three manipulation tasks:
TABLE II: Performance of our state-based policy using hand rewards at different stages. It follows the human hand trajectory extracted from videos to manipulate objects (R1), reach the pre-grasp configurations (R2) or combine both (R3).
@@
R1 $\times$ 0.063 0.15 0.37 0.00 0.00
R2 $\times$ 0.0031 0.053 0.93 0.32 1.00
R3 0.0031 0.024 0.95 0.83 1.00
Relocate task requires the robot to move an object to a target position. The task is successfully solved when the object reaches the target position within a certain threshold.
Pour task requires the robot to grasp a mug filled with particles and pour the particles into a container. The success rate is measured by the percentage of particles in the container.
Place inside task requires the robot to grasp a banana and place it into a mug. The success rate is measured by the percentage of the banana mesh in the mug.
Evaluation metrics. The success rate (${\rm SR}$) is the major metric for the above three tasks. For the relocate task, we adopt the 10cm threshold (${\rm SR_{10}}$) following [23, 28] to define the success rate. To measure the accuracy more precisely, we add a more rigorous threshold of 3cm and compute the success rate ${\rm SR_{3}}$. For the state-based policy, we additionally evaluate how well the learned physically plausible trajectory matches the reference trajectory in the video using four metrics [27]: ${\rm E_{o}}$ computes the average object position error between the current object trajectory and its reference trajectory over time; ${\rm E_{h}}$ computes the average finger tips position error between the current robot hand trajectory and its reference trajectory over time; ${\rm SR_{o}}$ reports the fraction of timesteps where ${\rm E_{o}}$ is below the 1cm threshold; ${\rm SR_{h}}$ computes the fraction of timesteps where ${\rm E_{h}}$ is lower than 5cm.
Implementation details. During RL training, we randomly set the object initial and target positions following [23] and also randomize the initial object orientation. During testing, we run 300 episodes with different initial configurations and report the average success rate. For the state-based policy, we use both robot and object states defined in Table I as inputs and predict the control command. The length of a single episode is 60, 80 and 100 for the relocate task, place inside task and pour task, respectively. We use PPO [66] to optimize the state-based policy. We set the learning rate and batch size to 1e-5 and 256, respectively. The RL training takes around 30 hours on a single V100 GPU and 32 CPUs for each video separately. Our visual policy only requires robot proprioceptive states (the left part of Table I) and point clouds as inputs. To train the visual policy, we rollout 500 successful trajectories from the state-based policy for each video, where trajectories differ in initial object positions, rotations and target positions. During rollouts, we render 3D scene point clouds from three depth cameras surrounding the table in the simulation platform. We train the visual policy with Adam optimizer using a learning rate of 1e-5 and a batch size of 64.It takes 8 hours on 4 A100 GPUs for our visual policy trained from 5 videos.
@@ IV-B Evaluation of the State-based Policy on the Relocate Task
We adopt the relocate task to evaluate the state-based policy. Specifically, we use one DexYCB video for each of the five object categories and train a state-based policy for each video separately. The initial object poses of the videos are illustrated in the first row of Figure 4.
TABLE III: Comparison with state-of-the-art methods on the relocate task for state-based policies. We optimize a state-based policy using a single human video for each object (S6, S7), while previous methods usually need about one hundred human videos for each object (S2-S4). * indicates that we re-train the method from [23] with 20 DexYCB videos for each object.
Methods Training algorithms Num. videos Mustard bottle Tomato can Sugar box Large clamp Mug Avg.
@@
S1 DexMV [23] TRPO [69] 0 0.06 - 0.67 - 0.00 - 0.51 - 0.49 - 0.35 -
S2 DexMV [23] SOIL [70] 97 0.33 - 0.98 - 0.67 - 0.89 - 0.71 - 0.72 -
S3 DexMV [23] GAIL+ [71] 97 0.06 - 0.66 - 0.00 - 0.52 - 0.53 - 0.50 -
S4 DexMV [23] DAPG [19] 97 0.93 - 1.00 - 0.00 - 1.00 - 1.00 - 0.79 -
S5 DexMV [23] DAPG [19] 20 1.00 0.63 1.00 0.56 0.09 0.00 0.08 0.00 0.02 0.00 0.44 0.24
S6 Ours / wo rot. PPO [66] 1 0.75 0.03 0.74 0.65 0.38 0.08 0.39 0.03 0.98 0.03 0.65 0.16
S7 Ours PPO [66] 1 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00
TABLE IV: Performance of separate visual policies for each object on the relocate task. The initial configurations in testing are derived from the first row of Figure 4 for each object. Our visual policies (V4-V6) only utilize one video in training. “Comp. PC” denotes complete point clouds sampled from object CAD models without occlusion.
Source Models Comp. PC Object states Visual inputs Num. Points Noisy Points Mustard bottle Tomato can Sugar box Large clamp Mug Avg.
V1 S4 512 $\times$ 1.00 0.67 1.00 0.50 0.14 0.00 1.00 0.77 1.00 0.58 0.83 0.50
V2 S4 $\times$ 512 $\times$ 1.00 0.62 1.00 0.48 0.14 0.00 1.00 0.76 0.99 0.55 0.83 0.48
V3 S4 $\times$ $\times$ 512 $\times$ 0.98 0.55 1.00 0.46 0.14 0.00 1.00 0.76 0.99 0.54 0.82 0.46
V4 S7 $\times$ $\times$ 512 $\times$ 1.00 0.89 1.00 0.86 1.00 0.96 1.00 0.91 1.00 1.00 1.00 0.92
V5 S7 $\times$ $\times$ 2048 $\times$ 1.00 0.97 1.00 0.80 1.00 0.98 1.00 0.93 1.00 1.00 1.00 0.94
V6 S7 $\times$ $\times$ 2048 0.97 0.77 0.97 0.57 1.00 0.88 1.00 0.89 1.00 1.00 0.99 0.82
Figure 5: Qualitative comparison of state-based policies using different rewards. R1 (w/o hand reward in pre-grasp) leads to unstable grasps. R2 (w/o hand reward in manipulation) results in unnatural hand actions. Our proposed policy R3 uses hand rewards at both stages and achieves the best performance.
@@ IV-B1 Ablation Studies
We first perform ablations to validate the importance of our hand rewards at the pre-grasp and the manipulation stages in state-based policy learning. Table II presents the averaged performance over the five videos for policies with different reward functions. The policy in R1 uses the proposed hand rewards (i.e., Equation 3) in the manipulation stage but follows the previous reward function [23, 28, 19] in the pre-grasp stage which is to minimize the distance between the robot hand and the object. It achieves poor performance due to the lack of guidance from human hand trajectory to approach the object. As shown in the first column of Figure 5, the policy is difficult to arrive at a plausible pre-grasp configuration to stably lift the object. The policy in R2 employs the proposed reward function in the pre-grasping stage (Equation 2) but follows PGDM [27] (i.e., the second and forth terms in Equation 3) in the object manipulation stage. The reward function in PGDM constrains the object motions to be close to the reference trajectory but does not add any constraints on the hand motions, which results in unnatural robot hand actions as shown in the second column of Figure 5. Our proposed policy in R3 utilizes hand rewards in both stages, which outperforms R1 and R2 on all the metrics and produces robust and realistic grasps as shown in the last column of Figure 5.
@@ IV-B2 State-of-the-art comparison
Table III presents the quantitative comparison of our state-based policy and state-of-the-art models [23]. S1 to S4 denote four model variants reported in [23]. As their training data is not released, we re-train their best model DAPG [19] (S4) using DexYCB videos and report the results in S5 for more fair comparison. We can see that the performance drops significantly for some objects, i.e., large clamp and mug. This indicates that their method requires sufficient amount of videos to learn the manipulation of objects with complex shapes. Our proposed policy in S7 significantly outperforms the previous methods despite only using one training video per object. Though S4 also performs well for mustard bottle, tomato soup can, large clamp and mug, it easily knocks down thin objects (e.g., sugar box) when the robot approaches the object and achieves poor performance. Our hand reward addresses this problem by mimicking the human pre-grasping trajectory. S6 is a variant of our model without rotation augmentation in training. The inferior performance of S6 to our final model S7 demonstrates the importance of augmentation to help improve the robustness of trained policies.
@@ IV-C1 Learning from a single video per object
We first explore training separate visual policies for each object with only one video as in the previous section IV-B. During testing, we use the same object configurations as in the section IV-B, which are derived from object poses in the first row of Figure 4 by randomly changing the initial object position and rotation around z-axis for each testing episodes. The results are presented in Table IV. From V1 to V3, we rollout trajectories from the S4 models in Table III  [^1] We use S4 models available at released codes of DexMV [23]. to train the visual policy. V1 has the same assumption as [28] which assumes known object states and complete object point clouds. V2 removes the dependency on complete point clouds but still requires object states. V3 further relaxes the assumption and only uses robot states together with scene point clouds. We can see that performance decreases when using no privileged object information (i.e., object states and complete object point clouds). From V4 to V6, we train the visual policy using rollouts from our state-based policies S7. Benefiting from high-quality trajectories generated by S7, our visual policy V4 can outperform V1 to V3 by a large margin without privileged information. To investigate how the density of point clouds affect the visual policy, we sample 2048 points in V5 at each timestep. This marginally improves the performance compared to V4. In V6 we evaluate the robustness of our visual policy by adding Gaussian noise with $\sigma$=3cm to 3D point clouds during training and testing. Our visual policy still functions reasonably under noisy point clouds observations.
TABLE V: Performance of separate policies for each object on the relocate task. The initial configurations in testing are derived from all three rows in Figure 4. We compare our visual policies using different number of training videos. “Mb.”, “Tc.”, “Sb.” and “Lc.” denote mustard bottle, tomato can, sugar box and large clamp respectively.
Num. videos Mb. Tc. Sb. Lc. Mug Avg.
S4 97 0.46 0.28 0.59 0.35 0.11 0.00 0.63 0.31 0.65 0.62 0.49 0.31
V5 1 0.33 0.32 0.41 0.25 0.33 0.32 0.43 0.31 0.58 0.57 0.42 0.35
V7 2 0.65 0.42 0.66 0.60 0.66 0.62 0.67 0.64 0.91 0.85 0.71 0.63
V8 3 0.96 0.66 0.95 0.82 0.89 0.63 0.98 0.92 0.98 0.97 0.95 0.80
TABLE VI: Performance of unified visual policies for all objects on the relocate task. The initial configurations in testing are derived from all three rows in Figure 4. We compare point cloud representations with different coordinate systems where “wo.”, “ta.” and “ha.” represent world, target and robot hand. “Mb.”, “Tc.”, “Sb.” and “Lc.” denote mustard bottle, tomato can, sugar box and large clamp respectively.
Coor. sys Mb. Tc. Sb. Lc. Mug Avg.
wo. ta. ha. ${\rm SR_{10}}$ ${\rm SR_{3}}$ ${\rm SR_{10}}$ ${\rm SR_{3}}$ ${\rm SR_{10}}$ ${\rm SR_{3}}$ ${\rm SR_{10}}$ ${\rm SR_{3}}$ ${\rm SR_{10}}$ ${\rm SR_{3}}$ ${\rm SR_{10}}$ ${\rm SR_{3}}$
V9 $\times$ $\times$ 0.77 0.32 0.73 0.44 0.80 0.29 0.87 0.40 0.79 0.58 0.79 0.41
V10 $\times$ 0.91 0.84 0.94 0.92 0.99 0.97 0.96 0.93 0.92 0.92 0.94 0.92
V11 0.94 0.86 0.99 0.98 0.98 0.93 0.99 0.97 0.96 0.95 0.97 0.94
IV-C2 Learning from multiple videos per object
We next introduce a more challenging testing scenario where we evaluate policies under three different initial object configurations for each object as shown in Figure 4. For example, the mustard bottle does not necessarily stand on the table, but can lie in different ways. Table V presents the results of different models. The state-based policy S4 [23] in Table III though employs privilege object information and multiple videos per object in training, still generalizes poorly to different initial object poses. Our visual policy V5 in Table IV also suffers in this challenging testing scenario due to the insufficient training videos. From V7 to V8, we gradually increase the number of training videos with diverse initial object configurations and demonstrate significant increase in performance.
Figure 6: Qualitative illustration of the performance of our visual policies on pour (a) and place inside (b) tasks.
@@ IV-C2 Learning from multiple videos per object
TABLE VII: Comparison with state-of-the-art methods on the pour and the place inside tasks for our policy models. Our approach uses a single human video for each task, while previous methods need more than 91 human videos.
@@
S1 DexMV [23] TRPO [69] 0 / 0 0.01 0.03
S2 DexMV [23] SOIL [70] 101 / 91 0.04 0.28
S3 DexMV [23] GAIL+ [71] 101 / 91 0.03 0.16
S4 DexMV [23] DAPG [19] 101 / 91 0.27 0.31
S7 Ours PPO [66] 1 / 1 0.97 0.68
V5 Ours Behavior cloning 1 / 1 0.97 0.68
IV-C3 Learning a unified multi-object visual policy from multiple videos
Finally, we train a single policy for all objects and configurations and report its performance in Table VI. We test the policy under three initial configurations for each object as the above paragraph. The visual policy V9 has the same architecture as the policy V8 in Table V, which uses 3D point clouds represented in the world coordinate frame. By comparing V9 and V8, we can see that learning a unified model is more challenging than training policies for each object separately. Therefore, we improve the point cloud representations in the visual policy. In V10, we transform the 3D point clouds to the target coordinate system. As result, the model becomes more aware of its target and improves the performance significantly, i.e., 79% to 94% under ${\rm SR_{10}}$. To further incorporate fine-grained hand-object interaction features, we additionally transform 3D points into six hand joint coordinate systems in V11 (i.e., five finger tips and the hand wrist). Benefiting from extracting rich hand-object interaction features, our final model achieves an average success rate of 97% under ${\rm SR_{10}}$ for the five objects under the challenging test scenario with three different configurations.

### v2

@@ III-C The Vision-based Policy
Though our state-based policy can work well for different manipulation tasks, it requires robot proprioceptive states and object states as inputs. However, reliably estimating object states is often not trivial in practice. To alleviate this issue, we propose a vision-based policy that only takes robot states and 3D scene point clouds as inputs. To generate training data for the visual policy, we rollout the optimized state-based policy and generate diverse physically plausible trajectories. During the rollout process, we compute the 3D point clouds ${{\rm PC_{w}}\in\mathbb{R}^{N\times 3}}$ from the depth camera, where $\rm PC_{w}$ is represented under the world coordinate system (i.e., the center of the table) and $N$ is the number of points.
Coordinate transformation. As shown in Figure 1, inspired by [75], we propose to transform $\rm PC_{w}$ into the desired target coordinate system $\rm PC_{t}$, which makes the model more aware of the target position for control predictions. To capture dense interaction features between the robot and the object, we additionally transform $\rm PC_{w}$ into different robot joint coordinate systems (i.e., palm and finger tips). Finally, we combine point clouds representations under different coordinate systems ${\rm PC}\in\mathbb{R}^{N\times 3(j+3)}$ and feed them into PointNet [76], where $j$ is the number of finger tips. Our visual policy predicts control commands based on extracted visual features and robot proprioceptive states.
Training. We train our visual policy using either the behavior cloning (BC) or the recently proposed 3D diffusion policy [77, 78] from generated physically plausible trajectories. The BC model directly takes transformed point clouds and robot states as inputs and predicts robot actions. The diffusion policy employs the extracted 3D features from PointNet [76] as the global condition for the denoising model and recovers actions from Gaussian noise. We train both models with $\ell 2$ loss between the predicted and ground-truth actions.
@@ IV Experiments
We conduct extensive experiments and evaluate state-based and visual policies learned by our approach on three manipulation tasks for five seen and ten unseen objects.
@@ IV-A Experimental Setting
Video dataset. The DexYCB [51] dataset contains human demonstration videos of hand-object interactions. Following [24], we focus on five objects including mustard bottle, tomato soup can, sugar box, large clamp and mug. For each object, we choose three videos which are most similar to those used in DexMV [^1] The videos in DexMV are not publicly released. in our experiments. Figure 2 presents motion retargeting results for these videos. Two evaluation protocols are used for benchmarking the performance of policies. Protocol #1 trains the policy for each object separately and uses initial object poses from the first row of Figure 2. For Protocol #2, we train a unified policy for all five objects and adopt three poses for each object shown in Figure 2. Furthermore, we evaluate the performance on ten unseen objects, namely master chef can, tuna fish can, pudding box, gelatin box, potted meat can, banana, pitcher base, bleach cleanser, wood block and foam brick.
Simulation environment. We follow previous works [16, 24, 34] by using the the Adroit robot hand and the MuJoCo simulator for fair comparison. However, this setup is less realistic, as the robotic hand is not attached to an arm and can move freely. To address this, we attach an Allegro robot hand to a UR5 arm, mirroring our actual hardware configuration. We use the SAPIEN [79] simulator for this setup, which simplifies creating this more realistic environment and allows to speed up simulation. Unless specified otherwise, the Adroit hand in MuJoCo is used for benchmarking against state-of-the-art methods, while the Allegro hand in SAPIEN is employed for training visual policies and performing ablation studies.
Manipulation tasks. We follow [24] to evaluate on three tasks. The first relocate task requires the robot to move an object to a target position. The pour task aims to grasp a mug filled with particles and pour the particles into a container. The success rate is measured by the percentage of particles in the container. The place inside task aims to grasp a banana and place it into a mug. The success rate is measured by the percentage of the banana mesh in the mug.
Figure 3: Qualitative comparison of state-based policies using different rewards for Protocol #1 and the Allegro hand. R1 (w/o hand reward in pre-grasp) leads to unstable grasps. R2 (w/o hand reward in manipulation) results in unnatural hand actions. Our proposed approach R3 uses hand rewards at both stages and achieves the best performance.
TABLE I: Performance of our state-based policy using different hand rewards (Allegro hand, Protocol #1). It follows the extracted human hand trajectory to manipulate objects (R1), reach the pre-grasp location (R2) or combine both (R3).
@@
R1 $\times$ 0.048 0.21 0.35 0.00 0.00
R2 $\times$ 0.0033 0.077 0.95 0.32 1.00
R3 0.0019 0.032 0.97 0.79 1.00
Evaluation metrics. The success rate (${\rm SR}$) is the major metric for the above three tasks. For the relocate task, we adopt the 10cm threshold (${\rm SR_{10}}$) following [24, 29] to define the success rate. To measure the accuracy more precisely, we add a more rigorous threshold of 3cm and compute the success rate ${\rm SR_{3}}$. For the state-based policy, we additionally evaluate how well the physically plausible trajectory matches the reference trajectory using four metrics [28]: ${\rm E_{o}}$ computes the average object position error between the current object trajectory and its reference trajectory over time; ${\rm E_{h}}$ computes the average finger tips position error between the current robot hand trajectory and its counterpart over time; ${\rm SR_{o}}$ reports the fraction of timesteps where ${\rm E_{o}}$ is below 1cm; ${\rm SR_{h}}$ computes the fraction of timesteps where ${\rm E_{h}}$ is lower than 5cm.
Implementation details. We use PPO [74] to optimize the state-based policy. The RL training takes around 2 hours on a single A100 GPU for each video. We rollout 100 successful trajectories from the state-based policy for each video, where trajectories differ in initial object positions, orientations and target positions. During rollouts, we render 3D scene point clouds from the depth camera. These data are used to train our visual policy, which takes robot joint positions and 3D point clouds as inputs. Training our visual policy on a single A100 GPU takes around 10 hours using behavior cloning and about 20 hours using the diffusion policy, both based on data from 15 videos. For testing, we run 300 episodes with different initial configurations and report the average success rate. The length of a single episode is 60, 80 and 100 for the relocate task, place inside task and pour task, respectively. The real robot takes around 1 minute to execute an episode.
@@ IV-B Evaluation of the State-based Policy on the Relocate Task
We evaluate the state-based policy on the relocate task using both the Adroit and Allegro hands. We follow Protocol #1 to train the state-based policy then evaluate each policy on the same object with novel placement of the object. Specifically, we randomly change the initial object position and rotation around z-axis for each testing episodes.
TABLE II: Comparison with state-of-the-art methods on relocate task for state-based policies. We optimize a state-based policy using a single human video for each object (S6, S7, S8) (Protocol #1), while previous methods usually need about 100 videos for each object (S2-S4). * indicates that we re-train the method from [24] with 20 DexYCB videos for each object.
Methods Robot Hands Training algorithms Num. videos Mustard bottle Tomato can Sugar box Large clamp Mug Avg.
@@
S1 DexMV [24] Adroit TRPO [80] 0 0.06 - 0.67 - 0.00 - 0.51 - 0.49 - 0.35 -
S2 DexMV [24] Adroit SOIL [81] 97 0.33 - 0.98 - 0.67 - 0.89 - 0.71 - 0.72 -
S3 DexMV [24] Adroit GAIL+ [82] 97 0.06 - 0.66 - 0.00 - 0.52 - 0.53 - 0.50 -
S4 DexMV [24] Adroit DAPG [16] 97 0.93 - 1.00 - 0.00 - 1.00 - 1.00 - 0.79 -
S5 DexMV [24] Adroit DAPG [16] 20 1.00 0.63 1.00 0.56 0.09 0.00 0.08 0.00 0.02 0.00 0.44 0.24
S6 Ours Adroit PPO [74] 1 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00
S7 Ours Allegro PPO [74] 1 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00
S8 Ours w/o rot. Allegro PPO [74] 1 0.85 0.85 1.00 0.99 0.99 0.97 0.98 0.95 0.97 0.94 0.96 0.94
TABLE III: Performance of visual policies for each object on relocate task with ${\rm SR_{3}}$ using Protocol #1 and Allegro hand.
Algorithms #Points Mb. Tc. Sb. Lc. Mug Avg.
V1 BC 512 0.99 0.82 0.82 0.78 0.90 0.86
V2 BC 2048 1.00 1.00 0.87 0.91 1.00 0.96
V3 Diffusion 2048 1.00 0.97 0.98 0.99 0.99 0.99
@@ IV-B1 Ablation Studies
We first perform ablations to validate the importance of our hand rewards at the pre-grasp and the manipulation stages in state-based policy learning. Table I presents the averaged performance over all objects for policies with different reward functions for the Allegro hand. The policy in R1 uses the proposed hand reward $R_{m}^{h}$ in the manipulation stage but follows the previous reward function [24, 29, 16] in the pre-grasp stage which is to minimize the distance between the robot and the object. It achieves poor performance due to the lack of guidance from the human hand trajectory to approach the object. As shown in the first column of Figure 3, the policy has difficulties in arriving at a plausible pre-grasp configuration to stably lift the object. The policy in R2 employs the proposed reward function $R_{p}$ in the pre-grasping stage but follows PGDM [28] (i.e., $R_{m}^{o}$ and $\mathds{1}_{\rm lift}$) in the manipulation stage. The reward function in PGDM constrains the object motions to be close to the reference trajectory but does not add any constraints on the hand motions, which results in unnatural robot actions as shown in the second column of Figure 3. Our proposed policy in R3 utilizes hand rewards in both stages, which outperforms R1 and R2 on all the metrics and produces more robust and realistic grasps as shown in the last column of Figure 3.
@@ IV-B2 State-of-the-art comparison
Table II presents the quantitative comparison of our state-based policy and state-of-the-art models [24]. S1 to S4 denote four model variants reported in [24] for the Adroit hand. As their training data is not released, we re-train their best model DAPG [16] (S4) using 20 DexYCB videos and report the results in S5 for fair comparison. We can see that the performance drops significantly for some objects, i.e., large clamp and mug. This suggests that their approach needs a large number of videos to effectively learn how to handle objects with complex shapes. Our proposed policy in S6 significantly outperforms all previous methods despite only using one training video per object. Though S4 also performs well for mustard bottle, tomato soup can, large clamp and mug, it easily knocks down thin objects (e.g., sugar box) when the robot approaches the object and achieves poor performance. Our hand reward addresses this problem by mimicking the human pre-grasping trajectory. As shown in S7, we achieve the same performance with the Allegro hand. S8 is a variant of our model S7 without rotation augmentation in the z-axis during training. The performance only drops insignificantly when tested for such rotations. This demonstrates the robustness of our method.
@@ IV-C1 Learning from a single video per object
We first investigate whether replacing the ground-truth object state in the state-based policy with visual point cloud inputs can affect the performance. For each video, we follow Protocol #1 to train a separate visual policy using the rollout data obtained from the state-based policy S7 in Table II, and evaluate under novel object placements as in the evaluation of the state-based policy. The results are presented in Table III. Compared with V1 that uses 512 points, V2 samples 2048 points and improves the overall performance, which shows the importance of point clouds densities. By comparing V2 and V3, the 3D diffusion policy [78] can achieve more robust performance than the plain behavior cloning algorithm. Our visual policies only perform slightly worse than our state-based policies.
TABLE IV: Performance of unified visual policies on relocate task under $\rm{SR}_{3}$ using Protocol #2 and Allegro hand. We compare the performance using different number of human videos and point clouds under different coordinate systems.
Algo. Vid. Coor. sys. Seen Unseen avg.
target hand Mb. Tc. Sb. Lc. Mug Avg.
V4 BC 1$\times$5 $\times$ $\times$ 0.33 0.33 0.33 0.30 0.42 0.34 0.31
V5 BC 2$\times$5 $\times$ $\times$ 0.55 0.57 0.36 0.56 0.66 0.54 0.33
V6 BC 3$\times$5 $\times$ $\times$ 0.69 0.80 0.83 0.86 0.88 0.81 0.38
V7 BC 3$\times$5 $\times$ 0.93 0.92 0.99 0.93 0.99 0.95 0.37
V8 BC 3$\times$5 0.93 0.97 0.97 0.98 1.00 0.97 0.41
V9 Diff. 3$\times$5 0.97 0.99 1.00 0.99 0.99 0.99 0.50
@@
TABLE V: Comparison with state-of-the-art methods on the pour and the place inside tasks using the Adroit hand. Our approach uses a single human video for each task, while previous methods need more than 91 human videos.
@@
L1 DexMV [24] TRPO [80] 0 / 0 0.01 0.03
L2 DexMV [24] SOIL [81] 101 / 91 0.04 0.28
L3 DexMV [24] GAIL+ [82] 101 / 91 0.03 0.16
L4 DexMV [24] DAPG [16] 101 / 91 0.27 0.31
L5 Ours PPO [74] 1 / 1 0.97 0.68
L6 Ours Diffusion 1 / 1 0.97 0.68
IV-C2 Learning a unified multi-object visual policy from multiple videos
Finally, we adopt Protocol #2 and train a single policy for all objects with the Allegro hand and report its performance in Table IV. We test the policy under three initial poses for each object shown in Figure 2. From V4 to V6, we gradually increase the number of videos for each object from one to three and observe a significant improvement in performance. Then, V7 additionally transforms the 3D point clouds into the target coordinate system. As a result, the model becomes more aware of its target and largely improves the average performance for seen objects from 81% to 95%. To further incorporate fine-grained hand-object interaction features, we further transform 3D points into hand joint coordinate systems in V8. Benefiting from extracting rich hand-object interaction features, V8 achieves even better performance than V7. Different from V8, V9 trains the visual model using 3D diffusion policy and achieves an average success rate of 99% under the challenging test scenario.
To investigate the generalization of our visual policies, we evaluate their performance using three different initial poses for ten unseen YCB objects. In Figure IV from V4 to V6, our model better generalizes to novel objects when using more videos. V8 incorporates fine-grained interaction features and enhances generalization. Compared with the behavior cloning, V9 significantly improves the unseen success rate from 41% to 50% by using the 3D diffusion policy, which demonstrates better robustness and generalization abilities.